# Some geometric series for Euler’s constant

### 07 May 2026 v2

@@ 4. The asymptotic behavior of $(e_{m})$
Let’s use $N$ rather than $n$ which has served earlier as summation index. Recall $\sum_{0\leq n<N}n^{m}=(m+1)^{-1}\bigl(B_{m+1}(N)-B_{m+1}(0)\bigr)$. The proof leading to (16) works identically with an arbitrary positive integer $N$ in the place of $2^{l}$, if $N$ is even. If $N\geq 3$ is odd, $N=2M+1$, $M\geq 1$, we have to be careful about how we split (the analog of) $q_{l}-p_{l}$ into two parts. Formerly we compared the first part with an integral from 0 to $\frac{1}{2}$ but now, using $0\leq n<\frac{1}{2}N$, i.e., $0\leq n\leq M$, we see that the rectangle basis extends a bit further to $\frac{1}{2}+\frac{1}{2N}$. We could improve that (for $m$ large) by bounding the contribution of $x=M/N$ pointwise by $\mathsf{e}^{-m/2}$ (as done in [6, Prop. 1] which handles general positive measures). But let us continue with $\int_{0}^{\frac{1}{2}+\frac{1}{2N}}\mathsf{e}^{m(x-1)}\mathrm{d}x<\mathsf{e}^{m/2N}m^{-1}\mathsf{e}^{-m/2}\leq\mathsf{e}^{m/6}m^{-1}\mathsf{e}^{-m/2}$, for the first part. For the second part, i.e. for the contributions of the points $x=N^{-1}n$, $M+1\leq n<N$, we can still use $x=\frac{1}{2}$ as reference point, there is no change, except that we could gain a little in the $\frac{m}{8}\mathsf{e}^{-m/2}$ term (as the weight of the half-interval is now not $1/2$, but $M/N$). Anyway, we end up with a slightly deteroriated final global upper bound $\mathsf{e}^{m/6}m^{-1}\mathsf{e}^{-m/2}+\frac{m}{8}\mathsf{e}^{-m/2}+18\mathsf{e}^{-2}m^{-2}$, which is $\leq C_{2}m^{-2}$ with $C_{2}=\max_{t\geq 0}(t\mathsf{e}^{-t/3}+\frac{t^{3}}{8}\mathsf{e}^{-t/2})+18\mathsf{e}^{-2}$. The maximum is again attained near $t=5$, it is found to be $\numprint{2.27}$ and $C_{2}<4.67$. ∎
As a corollary if we let both $m$ and $n$ go to infinity such that the ratio converges to some value $q$, we get (with $1$ on the right if $q=0$):
$\lim_{m,n\to\infty,mn^{-1}\to q}\frac{B_{m+1}(n)-B_{m+1}(0)}{n^{m+1}}=\frac{q}{\mathsf{e}^{q}-1}\,.$
This is indeed what one would expect from expressing the Bernoulli polynomials in terms of Bernoulli numbers and taking the limit term-wise, but (17) makes it easy (and gives an $O(m^{-1})$ error estimate). Reference [11], which uses the saddle-point method in the complex domain, provides full asymptotics for, among others, $B_{n}(nz+\frac{1}{2})$ as $n\to\infty$.
We now resume the study of the quantities $e_{m}$. Recall that we have set $L_{m}=\lfloor 2\log_{2}(m+1)\rfloor$ and already estimated
$0\leq\sum_{l>L_{m}}\bigl(1-(m+1)p_{l}\bigr)\leq\frac{1}{m+1},$
so, combining this with Equations (12) and (16):
$\sum_{0\leq l\leq L_{m}}\bigl(1-(m+1)q_{l}\bigr)\leq e_{m}\leq\sum_{0\leq l\leq L_{m}}\bigl(1-(m+1)q_{l}\bigr)+(L_{m}+1)(m+1)\frac{C_{1}}{m^{2}}+\frac{1}{m+1}\,.$
Let us recall Equation (14) defining $q_{l}$ and record our current result:
$e_{m}=\sum_{0\leq l\leq L_{m}}\bigl(1-(m+1)\frac{1-\mathsf{e}^{-m}}{2^{l}(\mathsf{e}^{m2^{-l}}-1)}\bigr)+O(\frac{\log m}{m}).$ (18)
This motivates the definition of auxiliary coefficients $f_{m}$ and $g_{m}$, for positive real $m$:
$f_{m}$ $=\sum_{l=0}^{\infty}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)$ (19)
$g_{m}$ $=-\sum_{l=-1}^{-\infty}\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}$ (20)
@@
Clearly the series defining $f_{m}$ converges uniformly for $m$ bounded. We bound trivially the terms in the series defining $g_{m}$::
@@
$0\leq-g_{m}\leq\frac{2m}{\mathsf{e}^{2m}-1}+\sum_{n=4}^{\infty}\mathsf{e}^{-nm/2}=O_{m\to\infty}(me^{-2m}).$
@@
Details are left to reader: $\sum_{l=0}^{\infty}(1-\frac{2^{-l}z}{\mathsf{e}^{2^{-l}z}-1})$ defines a meromorphic function in the entire complex plane, with poles at the elements of $2\pi i\mathbb{Z}\setminus\{0\}$, so in particular if we substitute $z=2^{t}$, we obtain an analytic function in the strip $|\Im t|<\frac{\pi}{2\log 2}$. Regarding $\sum_{k=1}^{\infty}\frac{2^{k}2^{t}}{\mathsf{e}^{2^{k}2^{t}}-1}$, it is absolutely and uniformly convergent if $a<\Re t<b$, $|\Im t|<\frac{\pi}{2}-\epsilon$, $0<\epsilon<\frac{\pi}{2}$, from bounding below the denominator and above the numerator. This proves that $\phi(t)$ is as stated. ∎
Returning to Equation (18), we get:
$e_{m}$ $=L_{m}+1-\frac{m+1}{m}(1-\mathsf{e}^{-m})\biggl(L_{m}+1-\sum_{0\leq l\leq L_{m}}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)\biggr)+O(\frac{\log m}{m})$
$=(1+O(\frac{1}{m}))\Bigl(f_{m}-\sum_{l>L_{m}}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)\Bigr)+O(\frac{\log m}{m}).$
For $l>L_{m}=\lfloor 2\log_{2}(m+1)\rfloor$, $2^{-l}m\leq m/(m+1)^{2}$, hence:
$\sum_{l>L_{m}}\bigl(1-\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\bigr)=O(\sum_{l>L_{m}}\frac{m^{2}}{4^{l}})=O(\frac{m^{2}}{4^{L_{m}}})=O(m^{-2}).$
This gives, for $m\to+\infty$:
$e_{m}=(1+O(\frac{1}{m}))\Bigl(\log_{2}(m)+\frac{1}{2}+\phi(\log_{2}(m))+O(m^{-2})+O(m\mathsf{e}^{-2m})\Bigr)+O(\frac{\log m}{m}).$
As $\phi$ is periodic, it is bounded and we have obtained the proof of the main estimate as given in Theorem 2:
@@ 5. A Fourier series
We have used the basic Riemann formula $\Gamma(s)\zeta(s)=\int_{0}^{\infty}x^{s-1}(\mathsf{e}^{x}-1)^{-1}\mathrm{d}x$ for $\Re s>1$. Using the usual bound $|\zeta(1+it)|=O_{t\to+\infty}(\log t)$ ([12, Thm. 3.5]), and $|\Gamma(1+it)|\sim\sqrt{2\pi|t|}\exp(-\frac{\pi}{2}|t|)$ from Stirling formula in the complex domain (or more expediently from the reflection formula $\Gamma(s)\Gamma(1-s)=\pi/\sin(\pi s)$), we recover the analyticity of $\phi(t)$ in the strip $|\Im t|<\pi/\log(4)$ which was stated in Proposition 4.
This completes the proof of Theorem 2.
@@ 6. Exponential generating function
The author did not find the next formula in standard references [1, 7, 8] although it is undoubtly well-known, and is related to the relation
@@
which, after canceling out the $1$ and multiplying by $z$ gives Equation (23). ∎
@@
Hence, if we let $E(t)$ denote the left-hand side of Equation (24), we see that $E(m)=(\mathsf{e}^{m}-1)f_{m}=(\mathsf{e}^{m}-1)\bigl(e_{m}+O(\frac{\log m}{m})\bigr)=\mathsf{e}^{m}e_{m}\bigl(1+O(m^{-1})\bigr)$ and thus $e_{m}=(1+O(m^{-1}))\mathsf{e}^{-m}E(m)=\mathsf{e}^{-m}E(m)+O(\frac{\log m}{m})$. This is very reminiscent of the general estimate from [6, Prop. 1] relating moments of some measures on the unit interval with their exponential generating function.
@@ 7. Concluding remarks
Comparison with Equation (21) shows that the above is exactly $\phi(\operatorname{Log}_{2}z)$. And indeed, the image of the right half-plane under $z\mapsto\operatorname{Log}_{2}z$ is the horizontal strip where we have defined $\phi$ as an nalytic function.

### 15 May 2026 v4

@@ 5. Approximations
The series defining $f_{m}$ has positive terms and is convergent. Lemmas 5 and 6 allow to bound $|e_{m}-f_{m}|=\Bigl|\sum_{l=0}^{\infty}\bigl(mq_{m}(2^{l})-(m+1)p_{m}(2^{l})\bigr)\Bigr|$, even after having moved the absolute values inside the summation. ∎
6. Interlude on another sequence
As part of [14, Ex. 6.3-34] the following is asked (we have translated to our notation). Let for $m\geq 2$
$a_{m}$ $=\sum_{l=1}^{\infty}\biggl(\sum_{1\leq k<2^{l}}(2^{-l}k)^{m-1}-\frac{2^{l}}{m}+\frac{1}{2}\biggr)$
$b_{m}$ $=\sum_{l=1}^{\infty}\biggl(\frac{1}{\mathsf{e}^{2^{-l}m}-1}-\frac{2^{l}}{m}+\frac{1}{2}\biggr).$
Show $a_{m}=b_{m}+O_{m\to\infty}(m^{-1})$. We note that the general term $a_{m}(l)$ in the $a_{m}$ series is $O(2^{-l})$. This follows from Euler-Maclaurin, or directly from using Bernoulli polynomials:
$a_{m}(l)=\frac{B_{m}(2^{l})-B_{m}(0)}{m2^{l(m-1)}}-\frac{2^{l}}{m}+\frac{1}{2}=\sum_{k=2}^{m-1}\binom{m}{k}m^{-k}B_{k}m^{k-1}(2^{-l})^{k-1}\,.$ (28)
We used $B_{0}=1$, $B_{1}=-\frac{1}{2}$ and isolated a term $m^{-k}$ to merge it with the binomial coefficient. Similarly to the proof of Lemma 6, we now take note of the inequality
$0\leq\frac{1}{k!}-\frac{\binom{m}{k}}{m^{k}}\leq\frac{1}{k!}\frac{k(k-1)}{2m}\,.$
Let $b_{m}(l)$ be the $l$th term in $b_{m}$. We see that we are going to obtain $\sum_{2^{l}>m}|b_{m}(l)-a_{m}(l)|=O(m^{-1})$ via the similar steps as in the proof of Lemma 6. It is important that in Equation (28), the smallest exponent of $m2^{-l}$ is $1>0$.
There remains to evaluate the contributions for $1\leq 2^{l}\leq m$. Using the notation from the previous section, it is a question of upper bounding
$\sum_{1\leq 2^{l}\leq m}\left|2^{l}p_{m-1}(2^{l})-2^{l}q_{m}(2^{l})\right|.$
This is at most:
$S_{1}+S_{2}=\sum_{1\leq 2^{l}\leq m}\left|2^{l}p_{m-1}(2^{l})-2^{l}q_{m-1}(2^{l})\right|+\sum_{1\leq 2^{l}\leq m}\left|2^{l}q_{m-1}(2^{l})-2^{l}q_{m}(2^{l})\right|\,.$
For $S_{1}$, we use $2^{l}\leq m$ and Lemma 4 to see that $S_{1}\leq mK_{4}(m-1)^{-2}=O(1/m)$. Regarding $S_{2}$, we estimate, for each positive integer $n$:
$(\mathsf{e}^{n^{-1}(m-1)}-1)^{-1}-(\mathsf{e}^{n^{-1}m}-1)^{-1}$ $=\frac{\mathsf{e}^{n^{-1}\xi}}{n(\mathsf{e}^{n^{-1}\xi}-1)^{2}}\quad(\xi\in[m-1,m])$
$\leq\frac{1}{n4\sinh^{2}((2n)^{-1}\xi)}\leq\frac{n}{\xi^{2}}\leq\frac{n}{(m-1)^{2}}\,.$
So $S_{2}\leq 2m/(m-1)^{2}=O(1/m)$.
Remark 7.
As we mentioned already, Theorem 4.1 from [17] (which probably implies the above $O(1/m)$ in view of its many parameters) has an invalid proof. It is apparent that our way of handling binomial coefficients should fix this issue, in general.
7. Asymptotics
@@
Clearly the series from Equation (27) defining $f_{m}$ converges uniformly for $m$ bounded (for now, $m$ is real positive). We define (at first, for positive real $m$):
$g_{m}=-\sum_{l=-1}^{-\infty}\frac{m2^{-l}}{\mathsf{e}^{m2^{-l}}-1}\,.$ (29)
We bound trivially the terms in the series defining $g_{m}$:
@@
$0\leq-g_{m}\leq\frac{2m}{\mathsf{e}^{2m}-1}+\sum_{n=4}^{\infty}\mathsf{e}^{-nm/2}=O_{m\to\infty}(me^{-2m}).$ (30)
@@
Details are left to reader: $\sum_{l=0}^{\infty}(1-\frac{2^{-l}z}{\mathsf{e}^{2^{-l}z}-1})$ defines a meromorphic function in the entire complex plane, with poles at the elements of $2\pi i\mathbb{Z}\setminus\{0\}$, so in particular if we substitute $z=2^{t}$, we obtain an analytic function in the strip $|\Im t|<\frac{\pi}{2\log 2}$. Regarding $\sum_{k=1}^{\infty}\frac{2^{k}2^{t}}{\mathsf{e}^{2^{k}2^{t}}-1}$, it is absolutely and uniformly convergent if $a<\Re t<b$, $|\Im t|<\frac{\pi}{2\log(2)}(1-\epsilon)$, $0<\epsilon<1$ (so that $\Re 2^{t}>2^{a}\sin\frac{\pi\epsilon}{2}$), from bounding below the denominator and above the numerator. This proves that $\phi(t)$ is as stated. ∎
Since $\phi(\log_{2}(m))=-\log_{2}(m)-\frac{1}{2}+f_{m}+g_{m}$, we now get from Proposition 3 and the bound (30):
@@ 8. A Fourier series
We have used the basic Riemann formula $\Gamma(s)\zeta(s)=\int_{0}^{\infty}x^{s-1}(\mathsf{e}^{x}-1)^{-1}\mathrm{d}x$ for $\Re s>1$. Using the usual bound $|\zeta(1+it)|=O_{t\to\pm\infty}(\log|t|)$ ([18, Thm. 3.5]), and $|\Gamma(1+it)|\sim\sqrt{2\pi|t|}\exp(-\frac{\pi}{2}|t|)$ from Stirling formula in the complex domain (or more expediently from the reflection formula $\Gamma(s)\Gamma(1-s)=\pi/\sin(\pi s)$), we recover the analyticity of $\phi(t)$ in the strip $|\Im t|<\pi/\log(4)$ which was stated in Proposition 4.
This is one among the Fourier series obtained by Szpankowski [17, Eq. (4.20)] and by Prodinger [16, Thm. 7]. The proof of Theorem 2 is complete.
@@ 9. Exponential generating function
For the proof we will use the Lemma 7 below, which is undoubtedly very well-know, but which we did not locate in standard references [1, 7, 8]. It also relates to the relation
@@
which, after canceling out the $1$ and multiplying by $z$ gives Equation (33). Another method starts from the identity $\sinh w=w\cosh(\frac{w}{2})\cosh(\frac{w}{4})\cosh(\frac{w}{8})\cdots$ and computes the logarithmic derivative at $w=\frac{z}{2}$. ∎
@@
Hence, if we let $E(t)$ denote the left-hand side of Equation (34), $E(m)=(\mathsf{e}^{m}-1)f_{m}=\mathsf{e}^{m}f_{m}+O(\log m)$, $f_{m}=\mathsf{e}^{-m}E(m)+O(\log(m)\mathsf{e}^{-m})$, $e_{m}=\mathsf{e}^{-m}E_{m}+O(m^{-1})$. This is very reminiscent of the estimate from [6, Prop. 1] relating moments of some measures on the unit interval with their exponential generating function. It seems that such relations are a well-known theme for the asymtotics of sequences arising from the theoretical analysis of search algorithms, see [14, 9].
@@ 10. Concluding remarks
Comparison with Equation (31) shows that the above is exactly $\phi(\operatorname{Log}_{2}z)$. And indeed, the image of the right half-plane under $z\mapsto\operatorname{Log}_{2}z$ is the horizontal strip where we have defined $\phi$ as an analytic function.
@@
Acknowledgements. Thanks to Jean-Paul Allouche for drawing the author attention to A372422 and to the work by Knuth and Prodinger.